# Cross-modal RAG: Sub-dimensional Retrieval-Augmented Text-to-Image Generation

### v2

@@ 4.1 Experiment Setup
  • Text-to-Image Retrieval Baselines: CLIP(ViT-L/14) radford2021learning is a widely adopted dual-encoder model pretrained on large-scale image-text pairs and remains the most commonly used baseline for T2I retrieval. SigLIP(ViT-SO400M/14@384) zhai2023sigmoid improves retrieval precision over CLIP by replacing the contrastive loss with a sigmoid-based loss. ViLLA varma2023villa is a large-scale vision-language pretraining model with multi-granularity objectives. GRACE (Structured ID) li2024generative is a recent generative cross-modal retrieval model. IRGen zhang2024irgen is a transformer-based image-text retriever. We report GRACE and IRGen results as cited from qutiger . GILL koh2023generating is a unified framework that combines generation and retrieval.
@@
  • Text-to-Image Generation Baselines: SDXL podell2023sdxl is a widely used high-quality T2I diffusion model. LaVIT jin2023unified is a vision-language model that supports T2I generation. RDM blattmann2022retrieval is a representative retrieval-augmented diffusion model. UniRAG sharifymoghaddam2024unirag is a recent retrieval-augmented vision-language model. GILL koh2023generating can perform both T2I-R and T2I-G.
@@
For T2I-R, we adopt the standard retrieval metric Recall at K(R@K, K=1, 5, and 10). For T2I-G, we evaluate the quality of generated images by computing the average pairwise cosine similarity of generated and ground-truth images with CLIP(ViT-L/14) radford2021learning , DINOv2(ViT-L/14) stein2023exposing , and SigLIP(ViT-SO400M/14@384) zhai2023sigmoid embeddings. We also employ style loss gatys2015neural to assess the artistic style transfer in the WikiArt dataset asahi417_wikiart_all .
@@
Dataset Construction We evaluate the text-to-image retrieval on the standard benchmark MS-COCO chen2015microsoft and Flickr30K young2014image test sets. As for the text-to-image generation, we evaluate the model’s image generation capabilities across different aspects and choose three datasets: artistic style transfer in the WikiArt asahi417_wikiart_all , fine-grained image generation in the CUB WahCUB_200_2011 , and long-tailed image generation in the ImageNet-LT liu2019large . For each genereation dataset, we select some test samples, and use the remaining as the retrieval database. More details in Appendix D.
@@
Implementation Details For T2I-R, our sub-dimensional dense retriever is composed of a pretrained CLIP vision encoder (ViT-L/14) and an adaptor. We train the sub-dimensional dense retriever on the COCO training set using the InfoNCE loss with a temperature of 0.07. The adaptor is optimized using the Adam optimizer with an initial learning rate of 5e-5, and a StepLR scheduler with a step size of 3 epochs and a decay factor of 0.6. For T2I-G, we use gpt-image-1 as our MLLM backbone and set $\beta=0.015$ based on Therorem 3.2. The experiments [^1] The code is available at https://github.com/mengdanzhu/Cross-modal-RAG. are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K@5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090.
@@ 4.2 Quantitative Evaluation of Text-to-Image Retrieval
Ours 80.78 97.00 99.16 97.50 100.00 100.00
We test our sub-dimensional dense retriever model with various types of T2I-R models. As shown in Tab. 1, our proposed sub-dimensional dense retriever achieves state-of-the-art performance across all metrics and outperforms all baselines by a substantial margin on both MS-COCO and Flickr30K datasets. On MS-COCO, our method achieves R@1 = 80.78%, R@5 = 97.00%, and R@10 = 99.16%, which are significantly higher than the best-performing baseline SigLIP. The relative improvements are 72% on R@1, 35% on R@5, and 20% on R@10, demonstrating our model’s superior capability in the text-to-image retrieval. Notably, it exhibits strong zero-shot T2I-R performance on Flickr30K, achieving near-perfect accuracy with R@1 = 97.50%, R@5 = 100.00%, and R@10 = 100.00%, and surpassing SigLIP’s R@1 = 82.20% by nearly 20%. These results confirm that our proposed sub-dimensional dense retriever significantly enhance fine-grained T2I-R compared to global embedding alignment such as CLIP and SigLIP, generative retrieval methods like GRACE and IRGen, region-based fine-grained match on ViLLA, and the retrieval and generation unified framework GILL.
@@ 4.3 Quantitative Evaluation of Text-to-Image Generation
Ours 0.746 0.604 0.744 0.019 0.764 0.600 0.744 0.815 0.761 0.812
We benchmark our Cross-modal RAG method against state-of-the-art text-to-image generation models, including diffusion-based (SDXL), autoregressive (LaVIT), retrieval-augmented (RDM, UniRAG), and retrieval and generation unified (GILL) baselines. The evaluation is conducted on three datasets that span different generation challenges: WikiArt (artistic style transfer), CUB (fine-grained), and ImageNet-LT (long-tailed). As shown in Tab. 2, Cross-modal RAG achieves the highest scores across all models and datasets. On WikiArt, our method achieves the best performance in CLIP, DINO, and SigLIP, along with the lowest style loss, indicating it can capture the particular artistic style specified in the retrieved images effectively. On CUB, Cross-modal RAG also performs strongly across all three metrics, because it can localize and leverage the specific visual details in the retrieved images to facilitate generation. On ImageNet-LT, Cross-modal RAG improves CLIP similarity by 22%, DINO by 89%, and SigLIP by 24% over the second-best SDXL. This indicates that our retrieval method can retrieve images that best match the query and only use the relevant entity for generation, which greatly benefits T2I-G in the long-tailed situation.
@@ 4.4 Qualitative Analysis
Figure 3: The retrieved pareto optimal images with their corresponding satisfied subqueries in (a) WikiArt, (b) CUB and (c) ImageNet-LT datasets and model generation results of Cross-modal RAG.
To qualitatively illustrate the effectiveness of our Cross-modal RAG model, we visualize some examples of our retrieved pareto optimal images with their corresponding satisfied subqueries and generated outputs across all datasets in Fig. 3. The satisfied subqueries of each retrieved Pareto-optimal image are non-overlapping, and each retrieved image is optimal with respect to the sub-dimensions it satisfies. Therefore, we can guarantee that the Pareto set $\mathcal{P}$ collectively covers images with all satisfied subqueries in the retrieval database $\mathcal{D}$. Moreover, since the model knows which subqueries are satisfied by each retrieved image, MLLM can be guided to condition on the relevant subquery-aligned parts of each retrieved image during generation. As shown Fig. 3(a), the model is capable of style transfer, learning the artistic style of a certain artist (e.g., Theodore Rousseau) while preserving the details corresponding to each subquery (e.g., road, forest, Fontainebleau). The model is also able to retrieve accurate fine-grained information and compose the entities in subqueries (e.g., black crown, black throat, yellow belly) to perform fine-grained image generation on the CUB dataset in Fig. 3(b). Moreover, the model is good at long-tailed or knowledge-intensive image generation. In Fig. 3(c), ImageNet-LT is a long-tailed distribution dataset with many rare entities (e.g., totem pole). Retrieving such correct images can help improve generation fidelity. Baseline models without retrieval capabilities tend to struggle in these scenarios. More comparisons of generated images with other baselines are provided in Appendix F.
@@ 4.5 Efficiency Analysis
As shown in Tab. 3, we compare the retrieval efficiency of CLIP (ViT-L/14) with our sub-dimensional dense and sparse retrievers on the COCO test set. Our sub-dimensional dense retriever is composed of a frozen CLIP encoder (ViT-L/14) and a lightweight adaptor. As reported in Table 1, the sub-dimensional dense retriever improves Recall@1 by +86.73% over CLIP on COCO. Despite the adaptor’s minimal overhead – only 0.01× CLIP’s GPU memory usage, 0.017× its number of parameters, and 0.5× its query latency – its performance gain is substantial. Our sub-dimensional sparse retriever is text-based and operates solely on the CPU, requiring no GPU memory consumption, no learnable parameters and achieving the lowest query latency. Our Cross-modal RAG method, a hybrid of our sub-dimensional sparse and dense retriever, can leverage the complementary strengths of both and achieve query latency that lies between the pure sparse and dense retrievers – closer to that of the sparse. These results show Cross-modal RAG’s efficiency and scalability for large-scale text-to-image retrieval tasks without compromising effectiveness.
@@ 5 Conclusion
We proposed Cross-modal RAG, a novel sub-dimensional retrieval-augmented text-to-image generation framework addressing domain-specific, fine-grained, and long-tailed image generation. Our method leverages a hybrid retrieval strategy combining sub-dimensional sparse filtering with dense retrieval to precisely align subqueries with visual elements, guiding a MLLM to generate coherent images on the subquery level. The Pareto-optimal image selection ensures the largest coverage of various aspects in the query. Extensive experiments demonstrated Cross-modal RAG’s superior performance over state-of-the-art baselines in T2I-R and T2I-G. The ablation study and efficiency analysis highlight the effectiveness of each component and efficiency in the Cross-modal RAG.
Appendix A The Prompt Example to Decompose Queries into Subqueries
@@ Appendix B Proof of the Time Complexity in Retrieval Efficiency
    We then discretize the simplex of subquery weights $\alpha$ into $K$ possible combinations. Each combination requires checking $\sum_{i}\alpha_{i}s_{i}\left(I_{j}\right)\text{ in }\mathcal{O}(\widetilde{N})\text{ time }$, thus $\mathcal{O}(K\times\widetilde{N})$ in total.
    Each adaptor pass handles both CLIP-based vision encoding (costing $T_{\text{clip}}$ and the adaptor’s own cross-attention (costing $T_{\text{adaptor}}$)). If we assume one pass per subquery set (of size $n$), the total cost is $\widetilde{N}\times n\times\left(T_{\text{clip }}+T_{\text{adaptor }}\right)$. Multiplied by $K$ weight vectors, this yields $\mathcal{O}\left(K\times\widetilde{N}\times n\times\left(T_{\text{clip }}+T_{\text{adaptor }}\right)\right)$.
@@ Appendix C Proof of Algorithm Optimality
We want $\Delta_{\text{sparse}}+\beta\Delta_{\text{dense}}>0$. In the worst case for $I_{a},\Delta_{\text{dense}}<0$, potentially as low as $-C_{\max}$. A sufficient condition for $I_{a}$ to stay preferred is
@@
We discretize the simplex $\left\{\alpha:\alpha_{i}\geq 0,\sum_{i}\alpha_{i}=1\right\}$. Because subqueries are strictly enumerated by $\alpha$, if an image satisfies a unique subquery set, it must appear as an $\arg\max\sum_{i}\alpha_{i}s_{i}\left(I_{j}\right)$ for some $\alpha$. Thus, no non-dominated $\mathbf{s}\left(I_{j}\right)$ is missed. $\Delta_{\text{dense}}$ can be further used to find an optimal image among those sharing the same $\mathbf{s}\left(I_{j}\right)$. Therefore, all Pareto-optimal solutions are obtained and $\mathbf{s}(I_{j})$ in the Pareto front $\mathcal{P}_{f}$ is unique.
@@ Appendix D Datasets
Regarding T2I-G, WikiArt dataset is a comprehensive collection of fine art images sourced from the WikiArt online encyclopedia. Our implementation is based on the version provided in [41]. To construct the test set, we compare artwork titles across different images and identify pairs that differ by at most three tokens. From each matched pair, we retain one sample, resulting in 2,619 distinct test examples. The query for each test sample is formatted as: <title> in the style of <artistName>. The retrieval database is composed of the remaining WikiArt images after excluding all test samples, ensuring no overlap between ground-truth and retrieval candidates, as shown in Tab. 5. The Caltech-UCSD Birds-200-2011 (CUB-200-2011) [44] is a widely used benchmark for fine-grained image classification and generation tasks. It contains 11,788 images across 200 bird species. We use the CUB dataset with 10 single-sentence visual descriptions per image collected by [46]. Similarly, to construct the test set, we compare captions across different images and identify pairs that differ by one token, resulting in 5,485 distinct test samples. The query for each test sample is formatted as: Draw a <speciesName>. <caption>. For each test sample, the retrieval candidates consist of all remaining images in the CUB dataset, excluding that test image. The ImageNet-LT dataset [45] is a long-tailed version of the original ImageNet dataset. It contains 1,000 classes with 5 images per class. We randomly choose one image from each class to construct the test samples. The retrieval database is composed of the remaining ImageNet-LT images after excluding all test samples. The query for each test sample is formatted as: A photo of <className>.

### v3

@@ 4.1 Experiment Setup
  • Text-to-Image Retrieval Baselines: CLIP (ViT-L/14) (Radford et al., 2021) is a dual-encoder model pretrained on large-scale image-text pairs and remains the most widely adopted baseline for T2I retrieval. SigLIP (ViT-SO400M/14@384) (Zhai et al., 2023) improves retrieval precision over CLIP by replacing the contrastive loss with a sigmoid-based loss. ViLLA (Varma et al., 2023) is a large-scale vision-language pretraining model with multi-granularity objectives. BLIP-2 is an efficient vision-language model (Li et al., 2023). We report BLIP-2 results as cited from (Ge et al., 2024). GILL (Koh et al., 2023) is a unified framework that combines retrieval and generation.
@@
  • Text-to-Image Generation Baselines: SDXL (Podell et al., 2023) is a widely used high-quality T2I diffusion model. FLUX.1-dev (Labs, 2024) is a T2I DiT-based model with high-fidelity outputs. LaVIT (Jin et al., 2023) is a vision-language model that supports T2I generation. RDM (Blattmann et al., 2022) is a representative retrieval-augmented diffusion model. UniRAG (Sharifymoghaddam et al., 2024) is a recent retrieval-augmented vision-language model. GILL (Koh et al., 2023) can perform both T2I-R and T2I-G. We use gpt-image-1 (OpenAI, 2025) and gemini-2.0-flash (Google, 2025) as our MLLM backbones, and compare them against their MLLM baselines without retrieval.
@@
For T2I-R, we adopt the standard retrieval metric Recall at K(R@K, K=1, 5, and 10). For T2I-G, we evaluate the quality of generated images by computing the average pairwise cosine similarity of generated and ground-truth images with CLIP(ViT-L/14) (Radford et al., 2021), DINOv2(ViT-L/14) (Stein et al., 2023), and SigLIP(ViT-SO400M/14@384) (Zhai et al., 2023) embeddings. We also employ style loss (Gatys et al., 2015) to assess artistic style transfer in WikiArt (Ushio, 2024).
@@
Dataset Construction We evaluate the text-to-image retrieval on the standard benchmark MS-COCO (Chen et al., 2015) and Flickr30K (Young et al., 2014) test sets. As for the text-to-image generation, we evaluate the model’s image generation capabilities across different aspects and choose three datasets: artistic style transfer in the WikiArt (Ushio, 2024), fine-grained image generation in the CUB (Wah et al., 2011), and long-tailed image generation in the ImageNet-LT (Liu et al., 2019). For each generation dataset, we select some test samples and use the remaining as the retrieval database. More details in Appendix D.
@@
Implementation Details For T2I-R, our sub-dimensional dense retriever is composed of a pretrained CLIP vision encoder (ViT-L/14) and an adaptor. We train the sub-dimensional dense retriever on the COCO training set using the InfoNCE loss with a temperature of 0.07. The adaptor is optimized using the Adam optimizer with an initial learning rate of 5e-5, and a StepLR scheduler with a step size of 3 epochs and a decay factor of 0.6. The finetuned CLIP (ViT-L/14) is finetuned on the same COCO training set. We set $\beta=0.015$ based on Therorem 3.2. The experiments [^1] The code is available at https://github.com/mengdanzhu/Cross-modal-RAG. are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K@5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090.
@@ 4.2 Quantitative Evaluation of Text-to-Image Retrieval
Ours 81.82 97.46 99.38 97.50 100.00 100.00
We test our sub-dimensional dense retriever model with various types of T2I-R models. As shown in Tab. 1, our proposed sub-dimensional dense retriever achieves state-of-the-art performance across all metrics and outperforms all baselines by a substantial margin on both MS-COCO and Flickr30K datasets. On MS-COCO, our method achieves R@1 = 81.82%, R@5 = 97.46%, and R@10 = 99.38%, which are significantly higher than the best-performing baseline BLIP-2. The relative improvements are 38.4% on R@1, 17.8% on R@5, and 11.2% on R@10, demonstrating our model’s superior capability in text-to-image retrieval. Notably, it exhibits strong zero-shot T2I-R performance on Flickr30K, achieving near-perfect accuracy with R@1 = 97.50%, R@5 = 100.00%, and R@10 = 100.00%, and surpassing CLIP’s R@1 = 77.40% by nearly 26%. These results confirm that our proposed sub-dimensional dense retriever significantly enhances fine-grained T2I-R compared to global embedding alignment such as CLIP, finetuned CLIP, SigLIP, and BLIP-2, region-based fine-grained match on ViLLA, and the retrieval and generation unified framework GILL.
@@ 4.3 Quantitative Evaluation of Text-to-Image Generation
gemini-2.0-flash 0.721 0.537 0.723 0.026 0.726 0.569 0.713 0.692 0.456 0.682
gpt-image-1 0.730 0.542 0.733 0.024 0.735 0.575 0.708 0.695 0.476 0.683
Ours (gemini-2.0-flash) 0.735 0.568 0.738 0.021 0.760 0.595 0.747 0.825 0.717 0.814
Ours (gpt-image-1) 0.746 0.604 0.744 0.019 0.764 0.600 0.744 0.815 0.761 0.812
We benchmark our Cross-modal RAG method against state-of-the-art text-to-image generation models, including diffusion-based (SDXL, FLUX), autoregressive (LaVIT, gemini-2.0-flash, gpt-image-1), retrieval-augmented (RDM, UniRAG), and retrieval and generation unified (GILL) baselines. The evaluation is conducted on three datasets that span different generation challenges: WikiArt (artistic style transfer), CUB (fine-grained), and ImageNet-LT (long-tailed). As shown in Tab. 2, our Cross-modal RAG achieves the highest scores across all models and datasets, outperforming both its MLLM backbone without retrieval and other text-to-image generation models. These results demonstrate that our retrieval module is highly generalizable across different MLLM backbones and more effective than existing retrieval-augmented methods. On WikiArt, our method achieves the best performance in CLIP, DINO, and SigLIP, along with the lowest style loss, indicating it can capture the particular artistic style specified in the retrieved images effectively. On CUB, Cross-modal RAG also performs strongly across all three metrics, because it can localize and leverage the specific visual details in the retrieved images to facilitate generation. On ImageNet-LT, our method can retrieve images that best match the query, which greatly benefits T2I-G in the long-tailed situation.
@@ 4.4 Qualitative Analysis
Figure 3: The retrieved pareto optimal images with their corresponding satisfied subqueries in (a) WikiArt, (b) CUB and (c) ImageNet-LT datasets and generation results of Cross-modal RAG.
To qualitatively illustrate the effectiveness of our Cross-modal RAG model, we visualize some examples of our retrieved pareto optimal images with their corresponding satisfied subqueries and generated outputs across all datasets in Fig. 3. The satisfied subqueries of each retrieved Pareto-optimal image are non-overlapping, and each retrieved image is optimal with respect to the sub-dimensions it satisfies. Therefore, we can guarantee that the Pareto set $\mathcal{P}$ collectively covers images with all satisfied subqueries in the retrieval database $\mathcal{D}$. Moreover, since the model knows which subqueries are satisfied by each retrieved image, MLLMs can be guided to condition on the relevant subquery-aligned parts of each retrieved image during generation. As shown Fig. 3(a), the model is capable of style transfer, learning the artistic style of a certain artist (e.g., Theodore Rousseau) while preserving the details corresponding to each subquery (e.g., road, forest, Fontainebleau). The model is also able to retrieve accurate fine-grained information and compose the entities in subqueries (e.g., black crown, black throat, yellow belly) to perform fine-grained image generation on the CUB dataset in Fig. 3(b). Moreover, the model is good at long-tailed or knowledge-intensive image generation. In Fig. 3(c), ImageNet-LT is a long-tailed distribution dataset with many rare entities (e.g., totem pole). Retrieving such correct images can help improve generation fidelity. Baseline models without retrieval capabilities tend to struggle in these scenarios. More comparisons of generated images with other baselines are provided in Appendix H.
@@ 4.5 Efficiency Analysis
As shown in Tab. 3, we compare the retrieval efficiency of CLIP (ViT-L/14) with our sub-dimensional dense and sparse retrievers on the COCO test set. Our sub-dimensional dense retriever is composed of a frozen CLIP encoder (ViT-L/14) and a lightweight adaptor. As reported in Table 1, the sub-dimensional dense retriever improves Recall@1 by +89.14% over CLIP on COCO. Despite the adaptor’s minimal overhead – only 0.01× CLIP’s GPU memory usage, 0.017× its number of parameters, and 0.5× its query latency – its performance gain is substantial. Our sub-dimensional sparse retriever is text-based and operates solely on the CPU, requiring no GPU memory consumption, no learnable parameters and achieving the lowest query latency. Our Cross-modal RAG method, a hybrid of our sub-dimensional sparse and dense retriever, can leverage the complementary strengths of both and achieve query latency that lies between the pure sparse and dense retrievers – closer to that of the sparse. These results show Cross-modal RAG’s efficiency and scalability for large-scale text-to-image retrieval tasks without compromising effectiveness.
@@ 5 Conclusion
We proposed Cross-modal RAG, a novel sub-dimensional text-to-image retrieval-augmented generation framework. By efficient and fine-grained T2I retrieval, it facilitates domain-specific, fine-grained, and long-tailed image generation. Our method leverages a hybrid retrieval strategy combining sub-dimensional sparse filtering with dense retrieval to precisely align subqueries with visual elements, guiding a MLLM to generate coherent images on the subquery level. The Pareto-optimal image selection ensures the largest coverage of various aspects in the query. Extensive experiments demonstrated Cross-modal RAG’s superior performance over state-of-the-art baselines in T2I-R and can also benefit T2I-G. The ablation study and efficiency analysis highlight the effectiveness of each component and efficiency in the Cross-modal RAG.
Appendix A Decomposing Queries $Q$ into Subqueries $q_{i}$
We decompose queries $Q$ into subqueries $q_{i}$ using the prompt template in Fig. 5 to obtain $t_{i}$. Furthermore, we remove the attribute modifiers in $q_{i}$ to derive $T_{i}$ with an LLM (e.g., GPT-4o mini). For example, if $t_{i}$ corresponds to the embedding of “orange cat”, then $T_{i}$ corresponds to the embedding of “cat”. In some cases where $q_{i}$ does not contain any attribute modifiers, $t_{i}$ and $T_{i}$ are identical.
@@ Appendix B Proof of the Time Complexity in Retrieval Efficiency
[…]
@@ Appendix C Proof of Algorithm Optimality
We want $\Delta_{\text{sparse}}+\beta\Delta_{\text{dense}}>0$. In the worst case for $I_{a},\Delta_{\text{dense}}<0$, potentially as low as $C_{\max}$. A sufficient condition for $I_{a}$ to stay preferred is
@@
We discretize the simplex $\left\{\alpha:\alpha_{i}\geq 0,\sum_{i}\alpha_{i}=1\right\}$. Because subqueries are strictly enumerated by $\alpha$, if an image satisfies a unique subquery set, it must appear as an $\arg\max\sum_{i}\alpha_{i}s_{i}\left(I_{j}\right)$ for some $\alpha$. Thus, no non-dominated $\mathbf{s}\left(I_{j}\right)$ is missed. $\Delta_{\text{dense}}$ can be further used to find an optimal image among those sharing the same $\mathbf{s}\left(I_{j}\right)$. Therefore, all Pareto‐optimal solutions are obtained and $\mathbf{s}(I_{j})$ in the Pareto front $\mathcal{P}_{f}$ is unique.
@@ Appendix D Datasets
Regarding T2I-G, WikiArt dataset is a comprehensive collection of fine art images sourced from the WikiArt online encyclopedia. Our implementation is based on the version provided in (Ushio, 2024). To construct the test set, we compare artwork titles across different images and identify pairs that differ by at most three tokens. From each matched pair, we retain one sample, resulting in 2,619 distinct test examples. The query for each test sample is formatted as: <title> in the style of <artistName>. The retrieval database is composed of the remaining WikiArt images after excluding all test samples, ensuring no overlap between ground-truth and retrieval candidates, as shown in Tab. 5. The Caltech-UCSD Birds-200-2011 (CUB-200-2011) (Wah et al., 2011) is a widely used benchmark for fine-grained image classification and generation tasks. It contains 11,788 images across 200 bird species. We use the CUB dataset with 10 single-sentence visual descriptions per image collected by (Reed et al., 2016). Similarly, to construct the test set, we compare captions across different images and identify pairs that differ by one token, resulting in 5,485 distinct test samples. The query for each test sample is formatted as: Draw a <speciesName>. <caption>. For each test sample, the retrieval candidates consist of all remaining images in the CUB dataset, excluding that test image. The ImageNet-LT dataset (Liu et al., 2019) is a long-tailed version of the original ImageNet dataset. It contains 1,000 classes with 5 images per class. We randomly choose one image from each class to construct the test samples. The retrieval database is composed of the remaining ImageNet-LT images after excluding all test samples. The query for each test sample is formatted as: A photo of <className>.